# Small but large: Single organic molecules as hybrid platforms for quantum technologies

### v1

@@ IV Open challenges
Several challenges need to be addressed both in theory and in experiment to unlock the potential of molecular host-guest systems discussed above. Here, we iterate on three major challenges that we identify from state-of-the-art understandings.
@@
First, advancing solid-state molecular quantum technology requires a deeper understanding of molecule–host interactions from first principles. Static interactions, such as the insertion of a single molecule into the host matrix, which influence ZPL frequencies and vibrational properties, can often be addressed through full or semi-classical simulations [132, 133, 134, 135]. However, current studies remain far from being quantitative due to the inherent inhomogeneity of local molecular environments [69]. Meanwhile, dynamical interactions including spectral diffusion, pure dephasing, non-radiative relaxations, and fluorescence blinking pose additional challenges, as they critically affect the optical properties of single molecules in the solid state by introducing decoherence [136]. These interactions not only hinder the exploitation of all molecular degrees of freedom but also constrain the choice of suitable host materials. A comprehensive understanding and characterization of such interactions are therefore key to the development of controllable and scalable molecular quantum technologies. However, modeling these interactions is particularly challenging due to the involvement of multiple degrees of freedom and their wide energy scales. Addressing this multi-scale problem requires a multidisciplinary approach that integrates tools from condensed matter theory and physical chemistry, including first-principles calculations and model systems. These theoretical efforts must also be supported by dedicated experiments to reveal the microscopic dynamics governing these systems. In the near term, these efforts may face significant challenges due to computational complexity, which could ultimately necessitate quantum computing resources. However, emerging machine learning technologies provide a promising avenue to tackle these complexities. For instance, deep learning strategies can be employed to extract insights from the vast datasets generated by joint experimental and theoretical efforts [137], enabling the prediction of molecular properties and their interactions with the environment. Beyond enhancing our understanding of microscopic dynamics, this line of research could also enable quantum control in molecular quantum technologies, opening new possibilities for engineering molecular systems with specific applications.
@@
A second major challenge lies in the material domain. Host materials for single molecules, which so far has been almost exclusively limited to organic materials apart from a recent demonstration with hexagonal boron nitride (hBN) [138], often lack physical and chemical stability. Photostable single molecules are typically found in crystalline materials with long-range order. This is normally produced through sublimation in vacuum or inert gas atmosphere. Single molecules in polymeric hosts such as polyethylene [27] and polymethyl methacrylate [139] have also been reported, however, with sub-optimal photostability. Recently developed fabrication methods including liquid phase reprecipitation [90, 140] and electrohydrodynamic nanoprinting [89] could produce crystals with sizes in the micrometer to sub-micrometer scale, and enable Fourier-limited emission [141]. However, these hosts are fragile due to their low melting temperature, making them prone to sublimation. Additionally, their solubility in common organic solvents makes them hardly compatible with standard lithographic processing steps. To address this challenge, efforts in experiment and theory should be joined to identify an essential and adequate set of criteria to benchmark host-guest combinations. The search for physically and chemically stable, large bandgap host materials within the vast chemical space could benefit from integrating computational molecular screening techniques. High-throughput methods and machine learning-based approaches can efficiently predict the stability, electronic properties, and compatibility of potential host materials, enabling rapid identification of promising candidates [142]. The recent discovery of significant line narrowing of single molecules on the surface of hBN [138] opens a promising prospect to extend the chemical space of host materials beyond organics.
@@
In addition to identifying new host materials, the deterministic integration of single molecules into nanoscale photonic and electronic devices is essential for achieving scalability in these technologies. A third challenge lies in the positioning of single molecules with the desired accuracy. Various approaches have been explored so far. For example, sparsely distributed nanocrystals could be identified through optical microscopy, and photonic structures can be fabricated directly around the selected molecule through direct laser writing [25]. Alternatively, sublimated flakes of the host material can be picked up using a fiber tip and transferred onto a desired location on a substrate using micro-manipulation [26]. Electrohydrodynamic nanoprinting has demonstrated the capability to place molecules on a substrate with lateral precision of approximately 100 nm [89]. Despite these advances, deterministic positioning of individual molecules with atomic-scale precision and assembly of multiple molecules into ordered arrays are still outstanding challenges. It is thus crucial to develop new methodologies for deterministically positioning molecules within the material. Notably, such deterministic positioning and ordering has been demonstrated in surface science experiments, for example, through tip manipulation in a scanning-tunneling microscope (STM) [143, 144]. However, the conditions of surface science experiments differ considerably with respect to those of quantum optics, from the required material substrate to vacuum conditions and to optical imaging techniques. Technical challenges need to be addressed to bridge the experimental settings of these two fields. Two-dimensional (2D) materials [138, 145] present a promising solution, potentially serving as substrates suitable for both STM manipulations and quantum optical measurements. For example, nanoscopic positioning and alignment of molecules on a 2D-material substrate can be performed on an STM. The substrate can be transferred for stacking of van-der-Waals hetero-structures and fabrication of quantum devices.
@@ V Discussions and conclusion
In addition to these paths, it is worth noting that the narrow ZPL of PAH molecules makes them highly sensitive probes of their local physical environment, such as electric fields and strain. Frequency tuning of the ZPL transition via the DC Stark effect has been routinely performed on single molecules. The demonstrated high electric field sensitivity of DBT molecules in 2,3‐dibromonaphthalene matrix theoretically enables the detection of single electrons in the vicinity of the molecule [146]. This capability open a up range of possibilities for interfacing with quantum electronic devices, such as superconducting circuits [147], electron-based qubit devices in semiconductors [148, 149] and 2D materials [150]. Furthermore, it will also be worth investigating molecules with a permanent dipole moment for providing higher charge sensitivity [151]. In addition, the large strain susceptibilities of these molecules could also be utilized to interface with other quantum systems. For example, molecular spin degrees of freedom [152] could potentially benefit from such phononic couplings, to achieve couplings through a phononic bus  [153]. Such phononic platforms could be realized with polymer host matrices feasible for nanostructuring, printing nanocrystals, or placing single molecules on conventional phononic structures [89, 38]. In a similar vein, pseudolocal vibrational modes could offer new ways to entangle molecules located at intermediate distances, bridging the electronic and phononic length scales and paving the way for the creation of collective many-body states.

### v2

@@ IV Open challenges
Several challenges need to be addressed both in theory and in experiments to unlock the potential of molecular host-guest systems discussed above. Here, we iterate on three major challenges that we identify from state-of-the-art understandings.
@@
First, advancing solid-state molecular quantum technology requires a deeper understanding of molecule–host interactions from first principles. Static interactions, such as the insertion of a single molecule into the host matrix, which influence ZPL frequencies and vibrational properties, can often be addressed through full or semi-classical simulations [156, 157, 158, 159]. However, current studies remain far from being quantitative due to the inherent inhomogeneity of local molecular environments [57]. Meanwhile, dynamical interactions including spectral diffusion, pure dephasing, non-radiative relaxations, and fluorescence blinking pose additional challenges, as they critically affect the optical properties of single molecules in the solid state by introducing decoherence [160]. These interactions not only hinder the utilization of all molecular degrees of freedom but also constrain the choice of suitable host materials. A comprehensive understanding and characterization of such interactions are therefore key to the development of controllable and scalable molecular quantum technologies. However, modelling these interactions is particularly challenging due to the involvement of multiple degrees of freedom and their wide energy scales. Addressing this multi-scale problem requires a multidisciplinary approach that integrates tools from condensed matter theory and physical chemistry, including first-principles calculations and model systems. These theoretical efforts must also be supported by dedicated experiments to reveal the microscopic dynamics governing these systems. In the near term, these efforts may face significant challenges due to computational complexity, which could ultimately necessitate quantum computing resources. However, emerging machine-learning technologies provide a promising avenue to tackle these complexities. For instance, deep learning strategies can be employed to extract insights from the vast datasets generated by joint experimental and theoretical efforts [161], enabling the prediction of molecular properties and their interactions with the environment. Beyond enhancing our understanding of microscopic dynamics, this line of research could also enable quantum control in molecular quantum technologies [162], opening new possibilities for engineering molecular systems with specific applications.
@@
A second major challenge lies in the material domain. Host materials for single molecules, which so far have been almost exclusively limited to organic materials apart from a recent demonstration with hexagonal boron nitride (hBN) [163], often lack physical and chemical stability. Photostable single molecules are typically found in crystalline materials with long-range order. This is normally produced through sublimation in vacuum or inert gas atmosphere. Single molecules in polymeric hosts such as polyethylene [31] and polymethyl methacrylate [164] have also been reported, however, with sub-optimal photostability. Recently developed fabrication methods including liquid phase reprecipitation [94, 165] and electrohydrodynamic nanoprinting [93] could produce crystals with sizes in the micrometer to sub-micrometer scale, and enable Fourier-limited emission [115]. However, these hosts are fragile due to their low melting temperature, making them prone to sublimation. Additionally, their solubility in common organic solvents makes them hardly compatible with standard lithographic processing steps. To address this challenge, efforts in experiment and theory should be joined to identify an essential and adequate set of criteria to benchmark host-guest combinations. The search for physically and chemically stable, large bandgap host materials within the vast chemical space could benefit from integrating computational molecular screening techniques. High-throughput methods and machine learning-based approaches can efficiently predict the stability, electronic properties, and compatibility of potential host materials, enabling identification of promising candidates [166]. The recent discovery of significant line narrowing of single molecules on the surface of hBN [163] opens a promising prospect to extend the chemical space of host materials beyond organics.
@@
In addition to identifying new host materials, the deterministic integration of single molecules into nanoscale photonic and electronic devices is essential for achieving scalability in these technologies. A third challenge lies in the positioning of single molecules with the desired accuracy. Various approaches have been explored so far. For example, sparsely distributed nanocrystals could be identified through optical microscopy, and photonic structures can be fabricated directly around the selected molecule through direct laser writing [29]. Alternatively, sublimated flakes of the host material can be picked up using a fiber tip and transferred onto a desired location on a substrate using micro-manipulation [30]. Electrohydrodynamic nanoprinting has demonstrated the capability to place molecules on a substrate with lateral precision of approximately 100 nm [93]. Despite these advances, deterministic positioning of individual molecules with atomic-scale precision and assembly of multiple molecules into ordered arrays are still outstanding challenges. It is thus crucial to develop new methodologies for deterministically positioning molecules within the material. Notably, such deterministic positioning and ordering have been demonstrated in surface science experiments, for example, through tip manipulation in a scanning-tunneling microscope (STM) [167, 168]. However, the conditions of surface science experiments differ considerably with respect to those of quantum optics, from the required material substrate to vacuum conditions and to optical imaging techniques. Technical challenges need to be addressed to bridge the experimental settings of these two fields. Two-dimensional (2D) materials [163, 169] present a promising solution, potentially serving as substrates suitable for both STM manipulations and quantum optical measurements. For example, nanoscopic positioning and alignment of molecules on a 2D-material substrate can be performed on an STM. The substrate can be transferred for stacking of van-der-Waals hetero-structures and fabrication of quantum devices.
@@ V Discussions and conclusion
In addition to these paths, it is worth noting that the narrow ZPL of PAH molecules makes them highly sensitive probes of their local physical environment, such as electric fields and strain. Frequency tuning of the ZPL transition via the DC Stark effect has been routinely performed on single molecules. The demonstrated high electric field sensitivity of DBT molecules in 2,3‐dibromonaphthalene matrix theoretically enables the detection of single electrons in the vicinity of the molecule [170]. This capability opens a range of possibilities for interfacing with quantum electronic devices, such as superconducting circuits [171], electron-based qubit devices in semiconductors [172, 173] and 2D materials [174]. Furthermore, it will also be worth investigating molecules with a permanent dipole moment for providing higher charge sensitivity [175]. In addition, the large strain susceptibilities of these molecules could also be utilized to interface with other quantum systems. For example, molecular spin degrees of freedom [152] could potentially benefit from such phononic couplings, to achieve couplings through a phononic bus  [176]. Such phononic platforms could be realized with polymer host matrices feasible for nanostructuring [29], printing nanocrystals [93], or placing single molecules on conventional phononic structures [40]. In a similar vein, pseudolocal vibrational modes could offer new ways to entangle molecules located at intermediate distances, bridging the electronic and phononic length scales and paving the way for the creation of collective many-body states.
@@ Acknowledgement
Data Availability
The data that support the findings of this article are openly available [177].